# Belief Bias Identification

### v2

@@ 2.2 Introducing multiple belief biases
This section shows how an extended set of biases can be introduced into the model.
@@ 2.2.1 Asymmetries between successes and failures & Preference-based biases
Equations (4) and (5) assume that deviations from the Bayesian agent in the likelihood are symmetric for successes and failures. That is, the agent over or under-reacts to “positive information” the same way she over or under-reacts to “negative” information. For a non-Bayesian agent this need not be the case. Consider instead that the agent weights successes and failures of the data generating process differently. Then her likelihood function would be $\tilde{L}(p|s_{1},\dots,s_{n})=\binom{n}{k}p^{\alpha k}(1-p)^{\beta(n-k)}$. In turn, her posterior would be beta distributed with shape parameters $\tilde{a}_{n},\tilde{b}_{n}$ such that:
@@
where $\alpha,\beta>0$. And $\alpha\neq\beta$ indicates asymmetric reactions to successes and failures in the DGP. Equations (6) and (7) are especially interesting if the agent has preferences over the state-space. Suppose this is the case; and suppose further that preferences are expressed by a utility function, which is continuous and monotonically increasing over $p$. Then, every success $k$ would be informative of a higher realisation of $p\in(0,1)$, i.e. a preferred state. Conversely, every failure $(n-k)$ is informative of a lower realisation of $p$. Therefore, successes and failures, can be interpreted as pieces of good and bad news respectively. This means that $\alpha>1$ (overreacting to positive information) or $\beta<1$ (underreacting to negative information) can be interpreted as optimism bias; while $\alpha<1$ or $\beta>1$ would indicate pessimism. Furthermore, $\alpha>\beta$ represents the good news effect, while $\alpha<\beta$ implies there is bad news effect. [^8] Asymmetries between successes and failures can also be present at the level of the prior. These distinctions are introduced in section 4, but left out of the main theoretical discussion. This is without loss of generality as the resulting equations are identical to equations (9) and (10) except for having separate $\rho$ and $\delta$ parameters for successes and failures. (see section 4.2)
@@ 2.2.2 Confirmation Bias
Confirmation bias is modeled as a positive correlation between overreaction to information signal $S$, and how confirming the signal is. The degree of confirmation of the signal is unrelated to the agent holding preferences [^9] A discussion about the distinction between motivated and unmotivated confirmation bias can be found in Zhenxun, (2023) over the state-space. In particular, the degree of confirmation $c$ is expressed as the area, in the density function of the prior, comprised between the expected value of the prior $E(\pi(p))$ and the mean of the information signal $k/n$. Formally:
@@ 2.2.3 Sequence-related biases and unrelated inference biases
Over(under)-reaction to information signals can also be driven by other reasons than the presence of preferences over the state-space. We consider two kind of alternative deviations: Sequence-related biases (i.e. hot hand fallacy and gambler’s fallacy) and over(under)-inference to information signals when preference over the state-space are not at stake.
@@
Sequence-related biases: Consider a partition of the signal space $S$. Namely, $S_{1}=\{s_{1}\dots s_{m}\}$ and $S_{2}=\{s_{m+1}\dots s_{n}\}$ where $n>m$. Suppose all of the $s_{i}\in S_{2}$ are either successes (i.e. $\sum_{n-(m+1)}^{n}s_{i}=n-m+1$, $s_{i}\in S$) or failures (i.e. $\sum_{n-(m+1)}^{n}s_{i}=0$, $s_{i}\in S$). In this context, the hot-hand fallacy is defined as inferring too much from information signals after observing a the last $n-m+1$ consecutive successes (i.e. $\alpha>1$) or failures (i.e. $\beta>1$). Conversely, gambler’s fallacy is defined as inference against the information signal after observing a the last $n-m+1$ consecutive successes (i.e. $\alpha<0$) or failures (i.e. $\beta<0$).
@@
Inference biases: Consider an almost equivalent state-space set, $\Omega_{NP}=(0,1)$ over which the agent form beliefs but where, contrary to $\Omega$, the agent does not hold any preference. So any element $\hat{p}$ of the set $\Omega_{NP}$, will have the characteristic that $u(\hat{p})=0$. Because this is the only difference between sets, only $\alpha$ and $\beta$ coefficients in equations (9) and (10) differ between forming beliefs over $\Omega$ or $\Omega_{NP}$. Therefore, when the agent forms beliefs over $\Omega_{NP}$, $\alpha$ and $\beta$ coefficients can be interpreted as over(under) inference for successes and failures, independent of preferences over the state-space.
@@
The biases in this subsection are only separably identified from one another when the agent faces multiple belief-elicitation decisions and observes different data generating processes coming from both $\Omega$ and $\Omega_{NP}$.
2.2.4 Under and overconfidence
@@
Confidence biases (overconfidence and underconfidence) are those which are strictly related to the variance of the agent’s posterior distribution in relation to the Bayesian variance. [^11] As an alternative to capture deviations in the second moment of the distribution, one could also use the concentration parameter of the Bayesian posterior ($\kappa_{n}=a_{n}+b_{n}$) to define under and overconfidence. This alternative definition has the advantage of isolating confidence biases within equations (9) and (10). (i.e. Underconfidence: $\alpha+\beta<2$, and Overconfidence: $\alpha+\beta>2$). However, it imposes harsh restrictions on $\alpha$ and $\beta$ parameters, making confidence biases harder to be detected. Overconfidence implies that the agent’s overall distribution of posterior beliefs is more dispersed than that of a Bayesian agent. Conversely, underconfidence yields a posterior distribution which is more dispersed over the values of $p$ than the distribution a Bayesian agent would have. Formally:
@@
where $\tilde{Var_{n}}$ is the variance of the agent’s posterior beta distribution, and $Var_{n}$ is the Bayesian variance. In equation (11), $\nu>1$ indicates overconfidence, while $\nu<1$ indicates underconfidence.
@@ 3 Experimental Design
The experiment tests the theoretical framework of section 2 both at the individual and population level. The experiment was conducted at the behavioral and Experimental Economics Laboratory (BEELab) at Maastricht University. 88 participants were recruited and, each of them had to make 30 belief-elicitation tasks. For each task, participants report their belief beta distributions twice, after observing two consecutive information signals. The average payment per participant was 15.9 euros. The experiment was pre-registered in October, 2023.121212See the preregistration at https://aspredicted.org/285v-b8cg.pdf The full instructions of the experiment can be found in section D of the appendix.
@@ 3.1 The belief-elicitation task
In each of the 30 tasks, participants observe a pool of 99 urns; each urn containing 100 balls. Each one of these urns contains a different distribution of red and blue balls. That is, Urn 1 has only one red ball and 99 blue, in Urn 2 there are only two red balls and 98 blue. This continues until Urn 99 where 99 balls are red and one is blue. Then, one of these urns is selected at random (i.e. selected from a uniform distribution), but the content of the urn is not revealed to the participants. Participants’ task is to guess the percentage of red balls in the selected urn (see Figure 2). In order to make their guesses, participants receive information signals by observing two sequences of balls, drawn with replacement. In the first sequence of draws, either one, two or three balls are extracted at random from the selected urn.
@@
After seeing this sequence participants are asked to report their belief distribution131313see section 3.2 for details about the elicitation of beta distributions. (Figure 3(a)) about the percentage of red balls in the selected urn. This first report, is taken as a prior. Once this prior distribution is elicited, a second sequence of draws is drawn from the same selected urn and shown to the subjects. In this case, either three, five or seven balls are extracted at random from the urn. After observing this second information signal, participants are asked to report their (posterior) belief distribution [^14] Both in the prior and posterior elicitations, participants need to update from a given default beta distribution. Before eliciting their prior, the default beta distribution they see is a uniform distribution (as this is the exogenously implemented “prior of the prior”). Accordingly, the default beta distribution they update before eliciting their posterior is their own prior report. (Figure 3(b)). Once this is done, a new urn from the pool of 99 urns is selected, with replacement, and this process is repeated thirty times. All of the subjects had an identical selection of urns and sequences of draws, but the order of the tasks was randomized between participants. All of the belief-elicitation tasks were incentiviced with a binarized scoring rule (more details in section 3.2 and appendix B)
To assess the role of motivated beliefs, fifteen of the thirty urns (placed at a random spots of the experiment) had an extra payment attached to it. I henceforth refer to these urns as dollar urns. In particular, when participants reported beliefs about these urns they receive a payment, in cents, equal to the (unknown) number of red balls in the selected urn. This means that participants should have a preference over the content of dollar urns.
@@ 3.2 Eliciting beta distributions
To elicit belief distributions I follow Gonzalez-Fernandez et al., (2024). This method is especially convenient if the elicited belief distribution follows a specific parametric form, as opposed to alternative non-parametric methods (Harrison et al.,, 2017; Crosetto and De Haan,, 2023).
In order to report their beliefs, participants were presented with a dynamic graphical interface that allows them to select their preferred beta distribution (See figures 3(a) and 3(b)). By moving two sliders, each associated with a different question, participants were able to select a specific beta distribution. Those questions were:
@@ 2. What is your uncertainty level about this percentage?
The first slider in figure, associated with question (1.), allows the subjects to manipulate the expected value of a beta distribution, while the second slider, associated with question (2.), allows the subjects to manipulate the standard deviation of such beta distribution.  [^15] The probability density function of any beta distribution can also be parameterized by its expected value and variance. Participants were shown a five-minute explanatory video on how to interpret the graph they selected and how to manipulate it. Subjects were specifically instructed to solve the task graphically, and they were also shown that the scale of the graph updates dynamically in order for the plot to be informative. Participants also had the option to enable the graph to update on a fixed scale if they desired to do so. Additionally, they answered related comprehension questions  [^16] In order to participate in the experiment subjects must have answered 3 out of 5 comprehension questions correctly. If any comprehension question was answered incorrectly, participants had a second chance to answer every question correctly and could test the software before starting to answer the relevant belief-elicitation tasks.
In order to avoid the possibility that participants reported bimodal beta distributions, the maximum range of the second slider (standard deviation) depends on the value of the first slider (expected value). The possibility of reporting bimodal beliefs was excluded for a number of reasons: First, it is challenging to interpret what a bimodal beta distribution means in the context of this experiment, secondly introducing a bimodal distribution could impair the understanding of the graphical interface by the subjects; and thirdly, a bimodal distribution could never be the Bayesian answer to the task, so reporting a Bayesian distribution is an option which is never disallowed to the participants.
@@
(a) Figure 3a: Example of prior belief elicitation after observing a first sequence of draws from a random urn showing “red, red”. This example shows a specific selection of a scaled beta distribution with an expected value of 56 and a standard deviation of 18.47
@@
(b) Figure 3b: Example of posterior belief elicitation after observing a second sequence of draws showing “blue, red, blue, red, blue”. The updated beta distribution is selected in this example shows an expected value of 51 and a standard deviation of 7.44
@@
To incentivize truthful reporting, a binarized scoring rule was implemented in every task. Specifically, I follow a similar method to the one suggested by Schlag et al., (2015), who propose different scoring rules to incentivize every moment of a probability distribution (see appendix B to understand the specific scoring rule). At the same time, I closely follow Danz et al., (2022) in not explicitly disclosing the exact scoring rule that is implemented. Instead, participants were told that they should always truthfully report their guess for the percentage red of balls, and their uncertainty level in the selected urn (i.e. the mean and standard deviation of the beta distribution) so as to maximize their expected payoff.
@@ 4 Baseline and Complete model regressions
This section presents the main analyses derived directly from the theoretical framework of section 2. In Section 4.1, the baseline model regression is outlined, introducing only a limited set of biases. Following that, Section 4.2 details the complete model (equations (14),(15) and (16)), encompassing all biases of the model. The decision to present these two models separately aims to facilitate comparison, investigating whether significant biases in the baseline model persist when accounting for multiple belief biases. Both models are executed and compared at both population and individual levels. In Section 4.3, a method is introduced to compare the relative importance of biases proven significant at the individual level.
@@ 4.1 Baseline model regressions
The baseline model, closely resembles equations (4) and (5). It tests for under(over)-inference and base rate neglect(overuse):
@@
In equations (12) and (13), successes and failures (variables $k$ and $(n-k)$), represent the number of red and blue balls observed in the second sequence of draws provided to participants. Variables $a_{0}$ and $b_{0}$ are the parameters of the beta distribution elicited by subjects after the first sequence of draws, while $\tilde{a}_{n}$ and $\tilde{b}_{n}$ are the parameters of the beta distribution elicited by subjects after observing the second sequence of draws. $\varepsilon_{a}$ and $\varepsilon_{b}$ are the error terms.
The parameter interpretation of equations (12) and (13) is akin to equations (9) and (10). Parameters $\gamma$ and $\delta$ similarly indicate under/over inference and base rate neglect/overuse respectively. However, equations (12) and (13) differ by acknowledging that these biases ($\gamma_{s}$,$\gamma_{f}$) and ($\delta_{s}$,$\delta_{f}$) may vary between successes and failures (i.e. realisations of red and blue balls). The presence of those biases is tested by comparing whether the estimated $\gamma$ and $\delta$ parameters are significantly different from their Bayesian values (i.e. $\gamma_{s}=\gamma_{f}=1$ and $\delta_{s}=\delta_{f}=1$).
Note that equations (12) and (13) do not have an intercept. This has a very straightforward theoretical justification. If all the independent variables in those equations take value 00 (i.e. $k=0,n=0,a_{0}=1,b_{0}=1$), both $\tilde{a}_{n}$ and $\tilde{b}_{n}$ take a value of $1$. This means that when one starts updating from a uniform prior distribution ($a_{0}=1,b_{0}=1$), and observes no information signals whatsoever ($k=0,n=0$), one must not update, i.e. remain at such uniform prior after updating ($\tilde{a}_{n}=1,\tilde{b}_{n}=1$). Having an intercept different than zero would imply that agents update in the absence of any kind of information.

### v4

@@ 2.2.1 Asymmetries between successes and failures & Preference-based biases
Equations (4) and (5) assume that deviations from the Bayesian agent in the likelihood are symmetric for successes and failures. That is, the agent over or under-reacts to “positive information” the same way she over or under-reacts to “negative” information. For a non-Bayesian agent this need not be the case. Consider instead that the agent weights successes and failures of the data generating process differently. Then her likelihood function would be $\tilde{L}(p|s_{1},\dots,s_{n})\propto p^{\alpha k}(1-p)^{\beta(n-k)}$. In turn, her posterior would be beta distributed with shape parameters $\tilde{a}_{n},\tilde{b}_{n}$ such that:
@@
where $\alpha\neq\beta$ indicates asymmetric reactions to successes and failures in the DGP. Equations (6) and (7) are especially interesting if the agent has preferences over the state-space. Suppose this is the case, and suppose further that preferences are expressed by a utility function, which is continuous and monotonically increasing over $p$. Then, every success $k$ would be informative of a higher realization of $p\in(0,1)$, i.e. a preferred state. Conversely, every failure (nk)(n-k) is informative of a lower realization of $p$. Therefore, successes and failures can be interpreted as pieces of good and bad news respectively. This means that $\alpha>1$ (overreacting to positive information) or $\beta<1$ (underreacting, or updating against the signal when confronted with negative information) can be interpreted as optimism bias; while $\alpha<1$ or $\beta>1$ would indicate pessimism. Furthermore, the model also captures motivated asymmetries: $\alpha>\beta$ represents the good news effect, while $\alpha<\beta$ implies there is bad news effect. [^8] Asymmetries between successes and failures can also be present at the level of the prior. These distinctions are introduced in section 4, but left out of the main theoretical discussion. This is without loss of generality as the resulting equations are identical to equations (9) and (10) except for having separate $\rho$ and $\delta$ parameters for successes and failures. (see section 4.2)
@@ 2.2.2 Confirmation Bias
Confirmation bias is modeled as a positive correlation between overreaction to information signal $S$, and how confirming the signal is. The degree of confirmation of the signal is unrelated to the agent holding preferences [^9] A discussion about the distinction between motivated and unmotivated confirmation bias can be found in Zhenxun (2023) over the state-space. In particular, the degree of confirmation $c$ is expressed as the area, in the density function of the prior, comprised between the expected value of the prior $E(\pi(p))$ and the mean of the information signal $k/n$. Formally:
@@ 2.2.3 Sequence-related biases and unrelated inference biases
Overreaction or underreaction to information signals can also be driven by other reasons different to the presence of preferences over the state-space. I distinguish between: sequence-related biases (e.g., the hot-hand fallacy and the gambler’s fallacy) and baseline inference. That is overreaction to information signals when preferences over the state-space are not at stake.
@@
Sequence-related biases: Consider a partition of the signal space $S$. Namely, $S_{1}=\{s_{1}\dots s_{m}\}$ and $S_{2}=\{s_{m+1}\dots s_{n}\}$ where $n>m$. Suppose all of the $s_{i}\in S_{2}$ are either successes (i.e. $\sum_{n-m}^{n}s_{i}=n-m$, $s_{i}\in S$) or failures (i.e. $\sum_{n-m}^{n}s_{i}=0$, $s_{i}\in S$). In this context, the hot-hand fallacy is defined as inferring too much from information signals after observing the last nmn-m consecutive successes (i.e. $\alpha>1$) or failures (i.e. $\beta>1$). Conversely, gambler’s fallacy is defined as inference against the information signal after observing a the last nmn-m consecutive successes (i.e. $\alpha<0$) or failures (i.e. $\beta<0$). Thus, sequence-related biases are modeled as a context-dependent component of inference distortions, rather than as a separate channel.
@@
Inference without preferences (baseline inference): Consider an almost equivalent state-space set, $\Omega_{NP}=(0,1)$ over which the agent form beliefs but where, contrary to $\Omega$, the agent does not hold any preference. So any element $\hat{p}$ of the set $\Omega_{NP}$, will have the characteristic that $u(\hat{p})=0$. Because this is the only difference between sets, only $\alpha$ and $\beta$ coefficients in equations (9) and (10) differ between forming beliefs over $\Omega$ or $\Omega_{NP}$. Therefore, when the agent forms beliefs over $\Omega_{NP}$, $\alpha$ and $\beta$ coefficients can be interpreted as over(under) inference for successes and failures, independently of preferences over the state-space.
@@
The biases in this subsection are only separably identified when the agent faces multiple belief-elicitation decisions that generate variation both in whether preferences are at stake (tasks in $\Omega$ versus $\Omega_{NP}$) and in whether the observed signal contains a streak.
2.2.4 Under and overprecision
@@
Precision biases (overprecision and underprecision) are those which are strictly related to the variance of the agent’s posterior distribution in relation to the Bayesian variance. [^11] As an alternative to capture deviations in the second moment of the distribution, one could also use the concentration parameter of the Bayesian posterior ($\kappa_{n}=a_{n}+b_{n}$) to define under and overprecision. This alternative definition has the advantage of isolating precision biases within equations (9) and (10). (i.e. Underprecision: $\alpha+\beta<2$, and Overprecision: $\alpha+\beta>2$). However, it imposes harsh restrictions on $\alpha$ and $\beta$ parameters, making precision biases harder to be detected. Underprecision implies that the agent’s overall distribution of posterior beliefs is more dispersed than that of a Bayesian agent. Conversely, overprecision yields a posterior distribution which is less dispersed over the values of $p$ than the distribution a Bayesian agent would have. That is:
@@
where $\tilde{Var_{n}}$ is the variance of the agent’s posterior beta distribution, and $Var_{n}$ is the Bayesian variance. In equation (11), $\nu<1$ indicates overprecision, while $\nu>1$ indicates underprecision.
Summary.
For ease of reference, Table 1 in Section 4.2 summarizes the mapping from parameter restrictions to the bias labels used throughout the paper.
@@ 3 Experimental Design
The experiment provides a setting in which the belief biases described in the theoretical framework of Section 2 can be identified at both the individual and population levels. It was conducted at the behavioral and Experimental Economics Laboratory (BEELab) at Maastricht University. A total of 88 participants were recruited, and each completed 30 belief-elicitation tasks. In each task, participants reported their belief distributions twice, after observing two consecutive information signals. The average payment per participant was 15.9 euros. The experiment was pre-registered in October 2023.121212See the preregistration at https://aspredicted.org/285v-b8cg.pdf The full instructions of the experiment can be found in Appendix F.
@@ 3.1 The belief-elicitation task
In each of the 30 tasks, participants observed a pool of 99 urns, each containing 100 balls. Each urn contained a different composition of red and blue balls. Thus, Urn 1 contained 1 red ball and 99 blue balls, Urn 2 contained 2 red balls and 98 blue balls, and so on, up to Urn 99, which contained 99 red balls and 1 blue ball. One of these urns was then selected at random from the pool, but its content was not revealed to participants. Their task was to guess the percentage of red balls in the selected urn (see Figure 2). To do so, participants received information signals by observing two sequences of balls drawn with replacement. In the first sequence, either one, two, or three balls were drawn at random from the selected urn.
@@
After observing this first sequence, participants were asked to report their belief distribution [^13] See Section 3.2 for details about the elicitation of beta distributions. (Figure 3(a)) about the percentage of red balls in the selected urn. This first report is treated as a prior. Once this prior distribution was elicited, a second sequence of draws from the same selected urn was shown to participants. In this case, either three, five, or seven balls were drawn at random from the urn. After observing this second information signal, participants were asked to report their posterior belief distribution [^14] Both in the prior and posterior reports, participants needed to update from a given default beta distribution. Before eliciting their prior, the default beta distribution shown was uniform (as this was the exogenously implemented “prior of the prior”). Accordingly, the default beta distribution shown before eliciting the posterior was the participant’s own prior report. (Figure 3(b)). Once this was done, a new urn from the pool of 99 urns was selected, with replacement, and the process was repeated 30 times. All subjects faced the same set of urns and sequences of draws, but the order of tasks was randomized across participants. All belief-elicitation tasks were incentivized using a binarized scoring rule (more details are provided in Section 3.2 and Appendix B).
To assess the role of motivated beliefs, fifteen of the thirty tasks, placed at random positions in the experiment, required participants to report beliefs about urns to which an additional payment was attached. I henceforth refer to these as dollar urns. In these tasks, participants received a payment in cents equal to the (unknown) number of red balls in the selected urn. Thus, participants had a monetary incentive to prefer urns with a higher proportion of red balls.
@@ 3.2 Eliciting beta distributions
To elicit belief distributions, I use the tool introduced in Gonzalez-Fernandez et al. (2025), which was validated there against non-parametric elicitation methods. This method is especially convenient in settings where beliefs are expected to follow a parametric form, while remaining closely connected to the broader distribution elicitation literature (Manski, 2004; Goldstein et al., 2008; Harrison et al., 2017; Crosetto and De Haan, 2023).
To report their beliefs, participants were presented with a dynamic graphical interface that allowed them to select their preferred beta distribution (see Figures 3(a) and 3(b)). By moving two sliders, each associated with a different question, participants were able to select a specific beta distribution. These questions were:
@@ 2. What is your uncertainty level about this percentage?
The first slider, associated with question (1), allows subjects to manipulate the expected value of a beta distribution, while the second slider, associated with question (2), allows them to manipulate its standard deviation. [^15] The probability density function of any beta distribution can also be parameterized by its expected value and variance. Participants were shown a five-minute explanatory video on how to interpret the graph they selected and how to manipulate it. Importantly, while participants adjusted sliders to report their beliefs, the relevant object for the analysis is the belief distribution they constructed, which was later mapped into beta-distribution parameters, rather than the individual slider positions themselves. Subjects were specifically instructed to solve the task graphically, and were also shown that the scale of the graph updates dynamically in order to keep the plot informative. Participants additionally had the option to fix the graph on a constant scale if they wished to do so. Before beginning the relevant belief-elicitation tasks, participants answered related comprehension questions [^16] In order to participate in the experiment, subjects had to answer at least 3 out of 5 comprehension questions correctly. If any comprehension question was answered incorrectly, participants were given a second chance to answer all questions correctly. and were allowed to test the software. For ease of interpretation, the interface did not allow subjects to report bimodal beta distributions. Bayesian reporting is nevertheless always feasible. [^17] Standard ex-post parametric fitting of the beta distribution imposes the same restriction (Engelberg et al., 2009).
More generally, because of the continuous nature of the variance slider, measurement error in uncertainty reports may affect bias identification. Appendix D.1 therefore applies a noise-injection procedure to assess the robustness of the results to measurement error in variance reports. This procedure imposes no restriction on the admissible variance range to rule out bimodality.
@@
(a) Figure 3a: Example of prior belief elicitation after observing a first sequence of draws from a random urn showing “red, red”. This example shows a specific selection of a scaled beta distribution with an expected value of 56 and a standard deviation of 18.47.
@@
(b) Figure 3b: Example of posterior belief elicitation after observing a second sequence of draws showing “blue, red, blue, red, blue”. The updated beta distribution selected in this example has an expected value of 51 and a standard deviation of 7.44.
@@
To incentivize truthful reporting, a binarized scoring rule was implemented in every task. Specifically, I follow a method similar to that suggested by Schlag and Van der Weele (2013), who propose scoring rules to incentivize different moments of a probability distribution (see Appendix B for the specific rule used here). At the same time, I follow Danz et al. (2022) in not explicitly disclosing the exact scoring rule to participants. Instead, participants were told that, in order to maximize their expected payoff, they should always truthfully report both their guess for the percentage of red balls in the selected urn and their uncertainty about that percentage (that is, the mean and standard deviation of the beta distribution they selected).
@@ 4 Baseline and Complete model regressions
This section presents the empirical specifications derived from the theoretical framework in Section 2. Section 4.1 introduces a baseline regression that captures only a limited set of belief biases. Section 4.2 then presents the complete specification (equations (14), (15) and (16)), which incorporates the full set of biases studied in the model. Estimating these two specifications side by side makes it possible to assess whether the bias patterns detected in the baseline model remain once richer bias channels are taken into account. Both models are estimated at the population and individual levels, since aggregate estimates may conceal substantial heterogeneity across subjects. Section 4.3 finally introduces a procedure to compare the relative importance of the biases at the individual level. Additionally, Appendix E reports a Grether-style reduced-form specification of the data in the traditional binary-state framework used in the belief-updating literature.
@@ 4.1 Baseline model regressions
The baseline model, closely resembles equations (4) and (5). It tests for under(over)-inference and base-rate neglect(overuse):
@@
In equations (12) and (13), successes and failures (variables $k$ and (nk)(n-k)), represent the number of red and blue balls observed in the second sequence of draws provided to participants. Variables $a_{0}$ and $b_{0}$ are the parameters of the beta distribution elicited by subjects after the first sequence of draws, while $\tilde{a}_{n}$ and $\tilde{b}_{n}$ are the parameters of the beta distribution elicited by subjects after observing the second sequence of draws. $\varepsilon_{a}$ and $\varepsilon_{b}$ are the error terms.
The parameter interpretation of equations (12) and (13) is akin to equations (9) and (10). Parameters $\gamma$ and $\delta$ similarly indicate under/over inference and base-rate neglect/overuse respectively. However, equations (12) and (13) differ by acknowledging that these biases ($\gamma_{s}$$\gamma_{f}$) and ($\delta_{s}$$\delta_{f}$) may vary between successes and failures (i.e. realizations of red and blue balls). The presence of those biases is tested by comparing whether the estimated $\gamma$ and $\delta$ parameters are significantly different from their Bayesian values (i.e. $\gamma_{s}=\gamma_{f}=1$ and $\delta_{s}=\delta_{f}=1$).
Note that equations (12) and (13) do not have an intercept. This has a very straightforward theoretical justification. If all the independent variables in those equations take value 0 (i.e. $k=0,n=0,a_{0}=1,b_{0}=1$), both $\tilde{a}_{n}$ and $\tilde{b}_{n}$ take a value of $1$. This means that when one starts updating from a uniform prior distribution ($a_{0}=1,b_{0}=1$), and observes no information signals whatsoever ($k=0,n=0$), one must not update, i.e. remain at such uniform prior after updating ($\tilde{a}_{n}=1,\tilde{b}_{n}=1$). Having an intercept different than zero would imply that agents update in the absence of any kind of information.